Under normal conditions (76% probability), Financing Plan A will produce a $30,000 higher return than Plan B. Under tight money conditions (24% probability), Plan A will produce $35,000 less than Plan B. What is the expected value of return?

A $100,800

B $114,400

C $14,400

D $13,600
Answer #1

Expected Value of Return = R1 * P1 + R2 * P2....Rn * Pn

Here in Given Problem

( Here assuming the return of Plan B = $ 1,00,000 Note : Becuase the return of Plan A is saying based on Plan B return )

         Then                 R1 = $ 1,30,000 { 1,00,000 + 30,000 }

                                   P1 = 0.76

                                   R2 = $ 65,000    { 1,00,000 - 35,000 }

                                  P2 = 0.24

Expected Value of return = 1,30,000 * 0.76 + 65,000 * 0.24

                                           = 98,800 + 15,600

                                           = 1,14,400

Option B is Correct.

Please comment, if any further assistance required.